# Eyes Wide Unshut: Unsupervised Mistake Detection in Egocentric Video by Detecting Unpredictable Gaze

### v2

@@ Abstract
In this paper, we address the challenge of unsupervised mistake detection in egocentric video through the analysis of gaze signals, a critical component for advancing user assistance in smart glasses. Traditional supervised methods, reliant on manually labeled mistakes, suffer from domain-dependence and scalability issues. This research introduces an unsupervised method for detecting mistakes in videos of human activities, overcoming the challenges of domain-specific requirements and the necessity for annotated data. By analyzing unusual gaze patterns that signal user disorientation during tasks, we propose a gaze completion model that forecasts eye gaze trajectories from incomplete inputs. The difference between the anticipated and observed gaze paths acts as an indicator for identifying errors. Our method is validated on the EPIC-Tent dataset, showing its superiority compared to current one-class supervised and unsupervised techniques.
Index Terms:
Gaze, Mistake, Unsupervised, Behaviour Understanding
I Introduction
Smart glasses have recently gained significant popularity, with various existing products available on the market capable of providing assistance to the user through Augmented Reality. In order to provide timely assistance, wearable devices should be able to identify moments in which the user makes mistakes or is confused and requires help [1]. If such instances are properly detected, the AI system can proactively support users by offering contextual information or suggestions on how to best carry out the task at hand [2]. Previous works tackled the problem of detecting mistakes from a fully supervised perspective, where mistake instances were labeled in egocentric video and a machine learning algorithm was trained to discriminate between video segments of correct action executions and incorrect ones [3, 2]. Such a fully supervised approach has two main downsides: 1) It is domain-dependent, hence requiring an accurate characterization of what a mistake is, depending on the context. For instance, a mistake in a kitchen scenario is different from a mistake in the assembly line; 2) It requires a sufficient number of manually labeled mistake instances, which is time consuming, requires expert knowledge and sometimes it is difficult to observe and record. Ideally, a wearable assistant should be able to infer when the behavioral patterns of the user deviate from the norm in order to determine if they need assistance in a scenario-independent setting, i.e., without making specific assumptions on how a mistake is defined and without requiring costly labels.
To overcome the limitations of previous approaches, we propose to detect mistakes in egocentric videos of human activity in an unsupervised way, by learning from unlabeled video.
@@ I Introduction
Figure 1: The proposed unsupervised mistake detection method. (a) We assume as input a video with a partial gaze trajectory on the initial part of the video. (b) A gaze completion model predicts a gaze trajectory for the remaining part of the video, conditioned on the input video and the partial trajectory. (c) A mistake is detected if the predicted trajectory is significantly different from the observed one, which suggests an anomalous attention pattern.
@@
We postulate that mistakes are characterized by abnormal behavior exhibited by the camera wearer, who may feel disoriented during the execution of the task at hand. As widely acknowledged in the literature, gaze patterns play an important role in the control of human activities [4], hence, we study how to analyze them as a means to perform mistake detection. We propose to train a model of human gaze behavior, in the form of a gaze predictor, and use it to predict likely eye gaze trajectories from video at inference time. Assuming that ground truth gaze trajectories are available as input, we frame gaze prediction as a novel “gaze completion” task in which a model takes as input a video and a partial gaze trajectory (Figure 1(a)) and is tasked to predict a likely continuation of the trajectory (Figure 1(b)). We expect video of correct action execution to represent normal user behavior, and hence to be characterized by predictable gaze patterns, while human behavior will be anomalous, and hence gaze will be unpredictable, when a mistake is made by the user. By comparing the predicted gaze with ground truth eye gaze trajectories estimated through a gaze tracker, we can determine whether a video segment is anomalous (Figure 1(c)).
Experiments on EPIC-Tent [5] show the effectiveness of the proposed method when compared with one-class anomaly detection methods [6, 7], and unsupervised mistake detection baselines.
In sum, the contributions of this work are as follows:
  • We investigate for the first time the problem of unsupervised mistake detection from egocentric video of human activity;
  • We propose a method to perform unsupervised mistake detection which leverages a gaze completion model to identify instances in which gaze patterns are unpredictable as a signal for human uncertainty;
  • We benchmark different approaches and baselines on the EPIC-Tent dataset, showing the effectiveness of the proposed method.
II Related Work
Our research is related to previous investigations on egocentric gaze estimation, use of gaze as an input signal in egocentric vision, mistake detection, and video anomaly detection.
II-A Egocentric gaze estimation
@@
Previous works considered the problem of estimating where a person’s gaze falls in the scene from the observation of egocentric video. Since gaze characterizes human behavior [4], models able to estimate human gaze from visual observations prove useful in augmented reality and assistive technologies downstream applications even when a gaze tracker is not available [8, 9]. The seminal work of [10] investigated the simultaneous recognition of daily activities and prediction of gaze location patterns with a probabilistic generative model, showing the advantages of modeling gaze for the understanding of human behavior. The same work also introduced GTEA Gaze and GTEA Gaze+, two of the first datasets of daily activities including egocentric video and human gaze trajectories. Following up on this seminal work, other investigations proposed approaches for gaze prediction from egocentric video incorporating egocentric cues [11], modeling task-dependent attention transition [12], leveraging vanishing point, manipulation point, hand regions [13], and proposing specific architectural changes [14]. Other datasets to study egocentric gaze estimation and its applications in a variety of scenarios have been proposed through the years, leading to further development in gaze estimation approaches. These include EGTEA Gaze+ [15], a subset of Ego4D [16], EPIC-Tent [17], HoloAssist [2], and IndustReal[18]. Lai et al. recently proposed GLC [19], a transformer-based model integrating diverse gaze cues which achieves state-of-the-art performance on the EGTEA Gaze+ and Ego4D benchmarks.
In this work, we leverage the state-of-the-art GLC model [19] as an off-the-shelf approach to learn a model of egocentric human attention. We show that such a model, if trained with proper data sampling and optimization objectives, is instrumental for unsupervised mistake detection in egocentric activity. Additionally, we introduce the novel problem of gaze completion, where partial gaze trajectories are observed by the predicted model, and show that such an approach achieves superior performance.
II-B Use of gaze in egocentric vision
@@
While many previous works focused on gaze estimation from video, few works investigated the use of gaze, estimated through a dedicated gaze tracker, as an input to support downstream egocentric vision applications. The seminal work of [8] exploited gaze for identifying task-relevant objects and their interaction modes from egocentric videos, achieving a 95% discovery rate across tasks. In [20], eye gaze movements are used to detect privacy-sensitive situations and preserve bystanders’ privacy. In [21], gaze is used to detect objects attended by the camera wearer. GazeGPT [22] utilizes eye tracking to assist large language models (LLMs) in classification tasks by discerning the user’s focal point within the world-facing camera view. GazeSAM [23] combines eye tracking with the Segment Anything Model for real-time object segmentation to allow radiologists to collect segmentation masks by looking at regions of interest during image diagnosis. GIMO [24] shows the advantages of incorporating eye gaze information for ego-centric human motion prediction. The work introduces a comprehensive dataset comprising body poses, scene scans, and ego-centric views with eye gaze, which are essential for accurately inferring human intent. TSM [25] is a hybrid text saliency model that merges cognitive reading models with human gaze supervision for Natural Language Processing (NLP). It showcases a robust correlation with human gaze across four datasets and suggests a joint modelling approach to incorporate TSM predictions into NLP tasks.
@@
In this paper, we illustrate the critical role of gaze in mistake detection. Our method involves comparing gaze trajectories predicted from visual data with ground truth gaze estimate through a gaze tracker of the wearable device to identify mistakes when the model’s prediction deviates from the ground truth.
II-C Mistake Detection in Egocentric Videos
@@
Mistakes naturally occur in human activities. The ability to automatically detect them from egocentric video can be beneficial for an AR assistant to offer support. A formal definition of this task was initially introduced in [3]. Identifying such mistakes usually entails modeling procedural knowledge [26], skill assessment [27], action segmentation [28] or detecting forgotten actions [29]. Notably, previous works tackled the task in a supervised fashion, training models to classify an action segment as “correct” or “mistake” on manually annotated instances [3, 2]. While this approach is feasible in closed-world scenario, it requires 1) a formal definition of what a mistake is, depending on the selected domain (e.g., kitchens vs the assembly line), 2) significant amounts of manually labeled data, which is expensive and requires expert knowledge. In this work, we tackle an unsupervised mistake detection task, in which models observe unlabeled video at training time and are tasked to output an anomaly score for given video segments at test time.
II-D Video Anomaly Detection
@@
Our research also relates to the problem of Video Anomaly Detection (VAD), which involves recognizing abnormal or anomalous events within videos [6, 30]. A line of video anomaly methods are based on one-class classification, in which models are trained only on normal videos. These approaches usually aim to learn a dictionary of normal features, which can either be obtained based on hand-crafted features [31] or deep autoencoder models [32]. Weakly-supervised learning methods, like Multiple Instance Learning (MIL), leverage both normal and anomalous videos during training, utilizing video-level labels to predict anomalies at the segment level [33, 34, 35]. Notably, anomaly detection in egocentric vision remains underexplored [36]. Similar to video anomaly detection, we aim to detect mistakes by determining video segments which deviate from statistics observed at training time. Differently from previous works in video anomaly detection, we ground our predictions in an egocentric gaze estimation model, which acts as a proxy for modeling human behavior, hence effectively mistake prediction detection when anomalous behavior is observed.
Figure 2: We consider three approaches to compare the ground truth with respect to the predicted trajectories in order to determine a mistake. (a) Euclidean distance between the two trajectories. (b) Dynamic Time Warping (DTW). (c) Average value of the ground truth trajectory at the predicted heatmaps. .
III Proposed Approach
III-A Mistake Detection Problem Setup
@@
The mistake detection task consists in highlighting those parts of the video in which the user is making a mistake during the execution of a given activity. In our setup, at each timestep $t$, a model $\Phi$ takes as input a video $V$ observed up to timestep $t$, $V_{1:t}$ and a 2D gaze trajectory $T_{1:t}$, where the $i-th$ element of the trajectory $T_{i}=(x_{i},y_{i})$ is a 2D gaze fixation, and have to return a score $s_{t}=\Phi(V_{1:t},T_{1:t})$ indicating whether a mistake is happening at the current time $t$. In this context, high $s_{t}$ scores indicate the occurrence of a mistake, while low $s_{t}$ scores indicate a correct action. We can hence see the mistake detection problem as a classification task, in which timesteps $t$ are classified as mistakes if $s_{t}>\phi$, where $\phi$ is a chosen threshold. We follow previous literature on anomaly detection [6, 7] and evaluate methods in a threshold-independent fashion by reporting the Receiver Operating Characteristics Area Under the Curve (ROC-AUC), where we consider “mistake” as the positive class [^1] In this context, a true positive is a mistake correctly classified as a mistake, a true negative is a correct execution correctly classified as a correct execution, a false positive is a correct execution wrongly classified as a mistake, and a false negative is a mistake wrongly classified as a correct execution.. For completeness, we also report the best $F_{1}$ score achieved considering the different threshold, as well as its related precision and recall values.
III-B Proposed Method
@@
Figure 3: The model inputs consist of $t$ observed RGB frames denoted as $V_{t-F:t}$ and the 2D gaze trajectory represented by $T_{t-F:t}$. The first half of the trajectory, $T_{t-F:t-F/2}$, highlighted in orange, is utilized in one of the two gaze fusion strategies. Following a peak finding operation, the latter part of the trajectory, $T_{t-F/2:t}$, is predicted.
III-B1 Gaze Completion Module
@@
Early and mid trajectory conditioning modules
Since GLC follows an encoder-decoder scheme, we experiment with two modules for injecting the input gaze trajectories into the model for the prediction of the output gaze trajectory.
@@
The early fusion module adds the heatmaps in $Q_{1:F/2}$ to the first $F/2$ frames of the input video, $V_{t-F:t-F/2}$ as an additional channel. The values of this channel are set to zero for the remaining frames. This acts as a form of soft conditioning aiming to include information about the input gaze trajectory in the computation. Note that, in order to incorporate information into the input gaze, the model needs to learn how to compute suitable gaze representations from the additional input channel.
@@
The mid fusion module processes the heatmaps $Q_{1:F/2}$ with a convolutional neural network encoder to produce a feature map, which is summed to the input of the decoder module of GLC. This module aims to encode gaze information in more direct way, with an additional convolutional neural network encoder components tasked to extract from the input gaze representations useful for GLC’s decoder.
Training Loss
Following [19], we consider gaze prediction as defining a probability distribution over the 2D image plane of each input frame. As in [19], we train the model by minimizing the sum of the Kullback–Leibler divergence between the predicted gaze maps $\hat{P}(i)$ and the ground truth ones $Q(i)$ at each frame $i$:
$L_{\text{KL}}(\hat{P}\parallel Q)=\sum_{i}\hat{P}(i)\log\left(\frac{\hat{P}(i)}{Q(i)}\right)$ (1)
III-B2 Scoring Function
@@
Our method predicts the mistake confidence score $s_{t}$ by comparing the predicted trajectory $\hat{T}$ with the ground truth one $T_{t-F:t-F/2}$ which is obtained by the gaze tracker of the wearable device. We consider three different ways to compare the two trajectories: Euclidean distance, Dynamic Time Warping and heatmap.
Euclidean Distance
@@
This method consists in using the Euclidean distance computed between corresponding points in each trajectory as the score $s_{t}$ (Fig. 2(a)):
$s_{t}=\sum_{i=0}^{F}\|T_{t-F+i}-\hat{T}_{i}\|$ (2)
where $\hat{T}_{i}=[\hat{x}_{t},\hat{y}_{t}]^{T}$ is predicted 2D gaze trajectory point at time step $t$.
Dynamic Time Warping
We then considered a more refined approach, utilizing Dynamic Time Warping (DTW) for comparison (Fig. 2(b)).
$DTW_{q}({T},{\hat{T}})=\min_{\pi\in\mathcal{A}({T},{\hat{T}})}\left(\sum_{(i,j)\in\pi}d(T_{i},\hat{T}_{j})^{q}\right)^{\frac{1}{q}}$ (3)
where $\pi$ is an alignment path, $\mathcal{A}({T},{\hat{T}})$ is the set of all admissible paths and $d$ is a distance measure.
Heatmap
@@
We consider a probabilistic approach which evaluates the likelihood of a ground truth eye fixation $T_{i}$ obtained by the device under the predicted heatmap $\hat{P}$, which can be computed as:
@@
$s_{t}=\sum_{i=0}^{F}P(T_{t-F+i}|\hat{P})$ (5)
Entropy
This method consists of calculating the mean entropy of all predicted heatmaps for a given trajectory $\hat{T_{i}}$. The entropy $H$ of a single heatmap $\hat{P}$ is given by:
$H(\hat{P})=-\sum_{i}\hat{p}_{i}\log_{2}(\hat{p}_{i})$ (6)
where $p_{i}$ denotes the normalized probability of the $i$-th bin in the heatmap.
IV Benchmark Experiments
In this section we discuss the datasets used for the experiments, the implementation details, and the experimental results.
IV-A Datasets
IV-A1 EPIC-Tent

### v3

# Eyes Wide Unshut: Unsupervised Mistake Detection in Egocentric Procedural Video by Detecting Unpredictable Gaze

@@ Abstract
In this paper, we address the challenge of unsupervised mistake detection in egocentric procedural video through the analysis of gaze signals. Traditional supervised mistake detection methods rely on manually labeled mistakes, and hence suffer from domain-dependence and scalability issues. We introduce an unsupervised method for detecting mistakes in videos of human activities, overcoming the challenges of domain-specific requirements and the need for annotated data. We postulate that, when a subject is making a mistake in the execution of a procedure, their attention patterns will deviate from normality. We hence propose to detect mistakes by comparing gaze trajectories predicted from input video with ground truth gaze signals collected through a gaze tracker. Since predicting gaze in video is characterized by high uncertainty, we propose a novel gaze completion task, which aims to predict gaze from visual observations and partial gaze trajectories. We further contribute a gaze completion approach based on a Gaze-Frame Correlation module to explicitly model the correlation between gaze information and each local visual tokens. Inconsistencies between the predicted and observed gaze trajectories act as an indicator for identifying mistakes. Experiments on the EPIC-Tent, HoloAssist and IndustReal datasets showcase the effectiveness of the proposed approach as compared to unsupervised and one-class techniques. Our method is ranked first on the HoloAssist Mistake Detection challenge.
@@
Figure 1: We aim to detect mistakes in egocentric procedural video by analyzing video and gaze fixations acquired through a gaze tracker, assuming that gaze fixations are unpredictable in the case of mistakes. Left: ground truth gaze trajectory of a correct and wrong execution of a “place water container” action, together with gaze fixation maps averaged across many correct and wrong executions of the action. Note the higher variability exhibited by wrong executions. Right: The proposed unsupervised mistake detection method. (a) We assume as input a video with a partial gaze trajectory on the initial part of the video. (b) A gaze completion model predicts a gaze trajectory for the remaining part of the video, conditioned on the input video and the partial trajectory. (c) A mistake is detected if the predicted trajectory is significantly different from the observed one, which suggests an anomalous attention pattern.
1 Introduction
Smart glasses have recently gained significant popularity, with various existing products available on the market capable of providing assistance to the user through Augmented Reality.
In order to provide timely assistance, wearable devices should be able to identify moments in which the user makes mistakes or is confused and requires help [6]. If such instances are properly detected, the AI system can proactively support users by offering contextual information or suggestions on how to best carry out the task at hand [36]. Previous works tackled the problem of detecting mistakes from a fully supervised perspective, where mistake instances were labeled in egocentric video and a machine learning algorithm was trained to discriminate between video segments of correct action executions and incorrect ones [28, 36]. Such a fully supervised approach has two main downsides: 1) It is domain-dependent, hence requiring an accurate characterization of what a mistake is, depending on the context. For instance, a mistake in a kitchen scenario is different from a mistake in the assembly line; 2) It requires to collect and label a sufficient number of mistake instances, which may be difficult to observe and record, involve time consuming procedures, and require expert knowledge. Ideally, a wearable assistant should be able to infer when the behavioral patterns of the user deviate from the norm in order to determine if they need assistance in a scenario-independent setting, i.e., without making specific assumptions on how a mistake is defined and without requiring costly labels.
To overcome limitations of previous approaches, we propose to detect mistakes in egocentric videos of human activity in an unsupervised way, learning from unlabeled video.
We postulate that mistakes are characterized by abnormal attention patterns exhibited by the camera wearer, who may deviate from a standard procedure or feel disoriented during the execution of the task at hand. Think of a user trying to operate the coffee machine but forgetting to add water first. As the user pushes the brew button, they may notice that no coffee is being brewed and shift their attention back and forth from the cup to the water tank to the led light signaling the correct functioning of the machine, a significant deviation from a standard “button $\to$ cup $\to$ button” attention pattern. This observation is grounded in previous literature showing that gaze patterns play an important role in the control of human activities [17].
@@ 1 Introduction
We propose to learn a model of “normal” attention patterns, in the form of a gaze predictor capable of outputting likely eye gaze trajectories from a video at inference time. Since gaze prediction can be governed by high uncertainty, depending on the user’s goals, we propose a novel “gaze completion” task in which a model takes as input a video and a partial gaze trajectory (Figure 1(a)) and is tasked to predict a likely continuation of the trajectory (Figure 1(b)). Gaze completion is tackled with a novel approach based on a Gaze-Frame Correlation module to explicitly model the correlation between gaze information and each local visual token. We expect videos of correct action executions to represent normal user behavior, and hence to be characterized by predictable gaze patterns, while human behavior will deviate from normality, and hence gaze will be unpredictable when a mistake is made by the user. We hence predict mistakes by comparing the predicted gaze with the ground truth eye gaze trajectory obtained through a gaze tracker(Figure 1(c)).
Experiments on EPIC-Tent [12], HoloAssist [36] and IndustReal[26], show the effectiveness of the proposed approach when compared to one-class anomaly detection methods [32, 7], and various unsupervised mistake detection baselines. In sum, the contributions of this work are as follows: 1) We investigate for the first time the problem of unsupervised mistake detection from egocentric video of human activity and provide an initial benchmark based on three different datasets, EPIC-Tent, HoloAssist, and IndustReal, and baseline results; 2) We define the novel “gaze completion” task consisting in predicting gaze trajectories from video and partial gaze inputs and introduce an approach to tackle this task based on a Gaze-Frame Correlation module; 3) We propose a method to perform unsupervised mistake detection which leverages a gaze completion model to identify instances in which gaze patterns are unpredictable as a signal for human uncertainty. Experiments show the effectiveness of the approach, while there is ample space for improvement in the newly proposed benchmark. We will publicly release the code and model checkpoints to support future research.
2 Related Work
@@
Egocentric gaze estimation Literature on gaze estimation from egocentric video is rich [4, 18, 11, 34, 1, 19, 16]. The seminal work of [4] investigated the simultaneous recognition of daily activities and prediction of gaze location patterns with a probabilistic generative model. Subsequent investigations proposed approaches for gaze prediction from egocentric video incorporating egocentric cues [18], modeling task-dependent attention transition [11], leveraging vanishing point, manipulation point, hand regions [34], and proposing specific architectural changes [1]. Other datasets to study egocentric gaze estimation and its applications in a variety of scenarios have been proposed through the years, leading to further development in gaze estimation approaches. These include EGTEA Gaze+ [19], a subset of Ego4D [10], EPIC-Tent [13], HoloAssist [36], and IndustReal[26]. Lai et al. recently proposed GLC [16], a transformer-based model integrating diverse gaze cues which achieves state-of-the-art performance on the EGTEA Gaze+ and Ego4D benchmarks.
We propose a novel gaze completion task and show its application to the problem of unsupervised mistake detection in egocentric video. We build on the state-of-the-art GLC model [16] to learn a model of egocentric human attention by tackling the proposed gaze completion task.
@@
Use of gaze in egocentric vision While many previous works focused on gaze estimation from video, few works investigated the use of gaze, estimated through a dedicated gaze tracker, as an input to support downstream egocentric vision applications. The seminal work of [2] exploited gaze for identifying task-relevant objects and their interaction modes from egocentric videos, achieving a 95% discovery rate across tasks. In [31], eye gaze movements are used to detect privacy-sensitive situations and preserve bystanders’ privacy. In [24], gaze is used to detect objects attended by the camera wearer. GazeGPT [15] utilizes eye tracking to assist large language models (LLMs) in classification tasks by discerning the user’s focal point within the world-facing camera view. GazeSAM [35] combines eye tracking with the Segment Anything Model [14] to allow radiologists to collect segmentation masks by looking at regions of interest during image diagnosis. GIMO [38] shows the advantages of incorporating eye gaze information for egocentric human motion prediction. TSM [29] is a hybrid text-saliency model merging cognitive reading models with human gaze for Natural Language Processing tasks.
@@
We show the effectiveness of gaze in mistake detection. Our method involves comparing gaze trajectories predicted from visual data with ground truth gaze estimated through a gaze tracker to identify mistakes when predictions deviates from the ground truth.
@@
Mistake Detection in Egocentric Videos Mistakes naturally occur in human activities. The ability to automatically detect them from egocentric video can be beneficial for an AR assistant to offer support. A formal definition of this task was initially introduced in [28]. Identifying mistakes usually entails modeling procedural knowledge [3, 27], skill assessment [8], action segmentation [9] or detecting forgotten actions [30]. Notably, previous works tackled the task in a supervised fashion, training models to classify an action segment as “correct” or “mistake” in manually annotated instances [28, 36]. While this approach is feasible in a closed-world scenario, it requires 1) a definition of what a mistake is, depending on the domain (e.g., kitchens vs the assembly line), 2) significant amounts of manually labeled data, which is expensive and requires expert knowledge.
In this work, we tackle an unsupervised mistake detection task, in which models observe unlabeled video at training time and are tasked to detect mistakes from video at test time. Our unsupervised scheme is possible through the analysis of gaze attention patterns, which provide a supervisory signal to create a joint video-gaze model of normal behavior.
@@
Video Anomaly Detection Our research also relates to the problem of Video Anomaly Detection (VAD), which involves recognizing abnormal or anomalous events within videos [32, 20]. A line of video anomaly methods are based on one-class classification, in which models are trained only on normal videos. These approaches usually aim to learn a dictionary of normal features, which can either be obtained based on hand-crafted features [21] or deep models [37]. Weakly-supervised learning methods, like Multiple Instance Learning (MIL), leverage both normal and anomalous videos during training, utilizing video-level labels to predict anomalies at the segment level [33, 22, 5]. Notably, anomaly detection in egocentric vision remains under-explored [23].
Similar to video anomaly detection, we aim to detect mistakes by determining video segments which deviate from statistics observed at training time. Differently from previous works in video anomaly detection, we ground our predictions in an egocentric gaze estimation model, which acts as a proxy for modeling normal human behavior, hence effectively achieving mistake prediction detection when anomalous behavior is observed. Moreover, we go beyond the one-class assumption and show that our method can also be used in an unsupervised setting in which unlabeled correct and mistake examples are included at training.
3 Proposed Approach
3.1 Mistake Detection Problem Setup
@@
The mistake detection task consists in highlighting those parts of the video in which the user is making a mistake during the execution of a given activity. In our setup, at each time-step $t$, a model $\Phi$ takes as input a video $V$ observed up to time-step $t$, $V_{1:t}$ and a 2D gaze trajectory $T_{1:t}$, where the $i-th$ element of the trajectory $T_{i}=(x_{i},y_{i})$ is a 2D gaze fixation in frame $V_{i}$. Given this input, the model has to return a score $s_{t}=\Phi(V_{1:t},T_{1:t})$ indicating whether a mistake is happening at the current time $t$. In this context, high $s_{t}$ scores indicate the occurrence of a mistake, while low $s_{t}$ scores indicate a correct action. We can hence see the mistake detection problem as a classification task, in which timesteps $t$ are classified as mistakes if $s_{t}>\phi$, where $\phi$ is a chosen threshold. We follow previous literature on anomaly detection [32, 7] and evaluate methods in a threshold-independent fashion by reporting the Receiver Operating Characteristics Area Under the Curve (ROC-AUC), where we consider “mistake” as the positive class [^1] In this context, a true positive is a mistake correctly classified as a mistake, a true negative is a correct execution correctly classified as a correct execution, a false positive is a correct execution wrongly classified as a mistake, and a false negative is a mistake wrongly classified as a correct execution.. For completeness, we also report the best $F_{1}$ score achieved considering the different thresholds, as well as its related precision and recall values.
3.2 Proposed Method
@@
Figure 2: The model takes as input $F$ RGB frames $V_{t-F:t}$ and a partial 2D gaze trajectory $T_{t-F:t-F/2}$ on the first $F/2$ input frames, and outputs a predicted gaze heatmap $\hat{P}$ from the input video, conditioned on the trajectory. The input trajectory is encoded as a spatio-temporal heatmap $Q$. Trajectory and RGB inputs are fused using a combination of three strategies, channel fusion, which adds gaze heatmaps as a separate channel (1), correlation fusion, using a dedicated gaze-visual correlation module (2) and sum fusion, summing $Q$ to a mid-level representation of the input (3). We follow the design of [16] and process our inputs with a transformer encoder-decoder architecture which outputs a predicted gaze heatmap $\hat{P}$, supervised via a standard Kullback–Leibler divergence loss using the ground truth unobserved trajectory $T_{t-F/2:t}$. The output trajectory $\hat{T}_{t-F/2:t}$ is recovered from $\hat{P}$ using a peak finding operation.
Gaze Completion Model Figure 2 illustrates the proposed gaze completion model. The model takes as input the video $V_{t-F:t}$ and the first half of the input gaze trajectory $T_{t-F:t-F/2}$ and predicts a gaze trajectory $\hat{T}_{t-F/2:t}$ aligned to the remaining part of the ground truth trajectory $T_{t-F/2:t}$. The goal of this model is to predict where the user is looking in the video, conditioned on an the initial trajectory. The conditioning allows to reduce the uncertainty on gaze predictions and give a prior into the intention of the user. We build on [16] and propose an encoder-decoder transformer-based architecture, experimenting with three approaches to condition gaze prediction on the input partial trajectory: channel fusion, correlation fusion, and sum fusion.
@@
Channel Fusion The channel fusion module (Figure 2(1)) adds the heatmaps in $Q_{1:F/2}$ to the first $F/2$ frames of the input video, $V_{t-F:t-F/2}$ as an additional channel. The values of this channel are set to zero for the remaining frames. This form of early fusion acts as a form of soft conditioning aiming to include information about the input gaze trajectory in the computation. Note that, in order to incorporate information into the input gaze, the model needs to learn how to compute suitable gaze representations from the additional input channel during training.
Correlation Fusion This approach (Figure 2(2)) aims to fuse visual and gaze tokens inspired by global-local fusion originally proposed in [16]. This is done in two stages. First, within the transformer encoder, where attention between all visual features and the gaze token is computed. In this stage, correlations between all visual tokens and across visual and gaze tokens are leveraged to obtain a strong representation. Second, within the dedicated gaze-visual correlation module. Here, attention is computed only between the gaze token and the visual token, which provides a much more focused mechanism to condition gaze prediction on the input trajectory. Note that this fusion mechanism operates both at the early (through the encoder) and mid (through the gaze-visual correlation module) levels.
@@
Sum Fusion The mid fusion module (Figure 2(3)) processes the input heatmaps $Q_{1:F/2}$ with a convolutional neural network encoder to produce a feature map which is summed to the input of the decoder transformer module. This module aims to inject gaze information at the mid level. Note that the CNN processing the input gaze heatmap is now tasked to extract from the input gaze representations useful for the decoder.
Figure 3: We consider three approaches to compare the ground truth with respect to the predicted trajectories in order to determine a mistake. (a) Entropy. (b) Euclidean distance between the two trajectories. (c) Dynamic Time Warping (DTW). (d) Average value of the ground truth trajectory at the predicted heatmaps.
@@
Scoring Function Our method predicts the mistake confidence score $s_{t}$ by comparing the predicted trajectory $\hat{T}_{t-F/2:t}$ with the ground truth one $T_{t-F/2:t}$ which is obtained by the gaze tracker of the wearable device. We explore four different ways to compare the two trajectories: Euclidean distance, Dynamic Time Warping (DTW), Entropy, and Heatmap. These are illustrated in Figure 3 and described in the following.
@@
Euclidean Distance This method consists in using the Euclidean distance computed between corresponding points in each trajectory as the score $s_{t}$ (Fig. 3(a)):
$s_{t}=\sum_{i=F/2+1}^{F}\|T_{t-F+i}-\hat{T}_{t-F+i}\|.$ (1)
Dynamic Time Warping This scoring function uses Dynamic Time Warping (DTW)[25] to measure the distance between trajectories (Fig. 3(b)):
$s_{t}=\min_{\pi\in\mathcal{A}({T},{\hat{T}})}\left(\sum_{(i,j)\in\pi}d(T_{i},\hat{T}_{j})^{q}\right)^{\frac{1}{q}}$ (2)
where $\pi$ is an alignment path (a sequence of pairs of points belonging to the two trajectories minimizing the sum of distances), $\mathcal{A}({T},{\hat{T}})$ is the set of all admissible paths (all paths following an increasing temporal ordering) and $d$ is a distance measure (we use the Euclidean distance). The reader is referred to this link [^3] https://pypi.org/project/fastdtw/ for more details.
@@
Heatmap Differently from previous functions, this approach explicitly considers the probability values predicted by the model at each location (Fig. 3(d)). Specifically, we evaluate the likelihood of a ground truth eye fixation $T_{i}$ obtained by the device under the predicted heatmap $\hat{P}$, which can be computed as:
@@
$s_{t}=\sum_{i=F/2+1}^{F}p(T_{t-F+i}|\hat{P}).$ (4)
Entropy This is the only method which does not require ground truth gaze for computation. We consider this measure as a way to check whether mistakes are systematically characterized by uncertain gaze predictions (Fig. 3(c)). In this case, the score $s_{t}$ is set as the mean entropy of all predicted heatmaps for a given trajectory $\hat{T_{i}}$. The entropy $H$ of a single heatmap $\hat{P}$ is given by:
$s_{t}=-\frac{1}{F/2}\sum_{i=F/2+1}^{F}\sum_{x,y}\hat{P}_{t-F+i}(x,y)\log_{2}(\hat{P}_{t-F+i}(x,y))$ (5)
where $P_{j}(x,y)$ is the value at coordinates $(x,y)$ of the heatmap $P_{j}$.
4 Experiments and Results
4.1 Datasets